leviisshort:

I NEED HELP ASAP A cylinder and a cone have the same diameter: 8 inches. The height of the cylinder is 5 inches. The height of the cone is 15 inches. Use π = 3.14. What is the relationship between the volume of this cylinder and this cone? Explain your answer by determining the volume of each and comparing them. Show all your work.

Aricuzyas:

@supie

AZ:

Do you know the formula for volume for a cylinder and a cone? @leviisshort

leviisshort:

wdym by that?

leviisshort:

all the info i have is in the question...

AZ:

Do you know how to calculate the volume of a cylinder? Or the volume of a cone? I'm guessing you don't so here: Volume for cylinder = \(\pi r^2h\) Volume for cone = \(\dfrac{1}{3} \pi r^2h\)

AZ:

what do you notice that is different between the two formulas?

leviisshort:

the 1/3

AZ:

Exactly!! So that means if we have the same radius (which is r) and the same height (which is h in the equation) that means a cone would have 1/3 the volume of that of a cylinder

AZ:

do you want to use the numbers that they gave you to prove that? `diameter: 8 inches. The height of the cylinder is 5 inches. The height of the cone is 15 inches.`

leviisshort:

yes please

AZ:

so they gave us the diameter how do we calculate the radius? remember, to find the radius just divide the diameter by 2

AZ:

so if the diameter is 8, what is the radius?

leviisshort:

4? i remember the radius being half of the diameter

leviisshort:

not 100% sure tho

AZ:

exactly!

AZ:

so now we have radius = r = 4 and they told us height = 5 for the cylinder and radius = 4 height = 15 for the cone

AZ:

so let's first focus on the cylinder volume = πr2h we said for the cylinder, r = 4 and h = 5 (and we can use 3.14 for \(\pi\) so put the numbers in volume = 3.14 * (4)^2 * 5

leviisshort:

251.2?

AZ:

Yes

leviisshort:

and id cut that into thirds or..?

leviisshort:

is that just my answer

AZ:

no so that was for your cylinder, let's take a look at the cone now, okay?

leviisshort:

ok

AZ:

the cone had a diameter of 8 which we said is the same thing as a radius of 4 but the cone has a height of 15

leviisshort:

so the cylinders volume is 251.2?

AZ:

so the formula for volume of a cone is Volume of cone = \(\dfrac{1}{3} \pi r^2 h\)

AZ:

yes

leviisshort:

so to find the volume for the cone i have to divide that by 3?

AZ:

If it had the same numbers, we could have but for the cone, they told us that the height is 15 and not 5 like the cylinder

AZ:

So let's plug in radius = 4 and height = 15 Volume = 1/3 * pi * (4)^2 * 15 Volume =

leviisshort:

i got 251.32? idk if thats right it sounds weird

AZ:

It is!

AZ:

So remember how we said that if BOTH the radius and height were the same the volume of a cone would be 1/3 of that of a cylinder

leviisshort:

ya

AZ:

but this time, what they did was they gave you the same radius BUT the height of the CONE was 3 times the height of the cylinder

AZ:

so the 1/3 cancelled out with the 3 times the height to give you the same volume for both

AZ:

does that make sense?

leviisshort:

mhm

leviisshort:

thank u :D

AZ:

No problem!! :D